OpenStudy (qaisersatti):

anyone who has started Homework 12....

OpenStudy (anonymous):

At the moment) H12P1 A) Just V/i b) first find R2 = VS / i second the anwser is R = R2 - R1 (R1 is the awnser in a) c) I'm Trying LAB12 R1 (in Ohms): 1k C1 (in Farads): 1.59n R2 (in Ohms): 100k C2 (in Farads): 1.59n

OpenStudy (anonymous):

Jose.oliveira2 R1 (in Ohms): 1k C1 (in Farads): 1.59n R2 (in Ohms): 100k C2 (in Farads): 1.59n These don't work can u upload the pic of ur op-amp

OpenStudy (anonymous):

R1=10 R2=2490 C1=0.0000001 C2=0.0000001

OpenStudy (anonymous):

H12P3 a) ? b) ? c) 20000 d) 0.82u e) 390p f) ? g) ? h) 0.82u i) 390p j) ?

OpenStudy (qaisersatti):

Our specification is that we want there to be two break frequencies for this filter: 10Hz and 17kHz. We also want midband frequencies, such as signals around 1kHz, to be amplified by about a factor of 2. We have available a 1.8μF and a 220pF capacitor. We also have a variety of resistors between 100Ω and 100kΩ. However, our op amp really likes RS=20kΩ. Your job is to fill in the remaining details of this design. Q No 3 answers

OpenStudy (anonymous):

HWP3 - first 5 filter: 10Hz and 19kHz. signals around 1kHz, to be amplified by about a factor of 2. 0.68μF and a 390pF capacitor. resistors between 100Ω and 100kΩ. RS=20kΩ. 1) First find C1=0.68*10^-6 = 6.8e-7 (answer of part 4) Then R1*C1=1/(2*pi*f) R1*C1=1/(2*pi*10) so R1= 23417 2)First find C2=390*10^-12 = 3.9e-10 (answer of part 5) R2*C2=1/(2*pi*19*10^3) R2= 21489 3) Rs = Rf = 20kΩ = 2000 Ω

OpenStudy (anonymous):

Remaining answers HWP3 part 6 : R3= R1 part 7 : R4= R2 part 8: c3=c1 part 9 : c4= c2 part 9 :Midband gain = -­(R4/R3)

OpenStudy (anonymous):

some one asked me the value of pi. its always 3.14 buddy!

OpenStudy (anonymous):

H12 P1 c) Q: v=5v A=100, load resistance =500 Solve two equations to find value of vx vx/R=(vOut–vx)/500,with R=answer from Part1. vx=5–vOut/100 Solving two equations we get Vx Current = Vx / R

OpenStudy (anonymous):

@machu I did not undesrtand the c part of H12P1. I have these values: i=6mA A=100 R=833.333333333 VS=20V RL=500Ω

OpenStudy (qaisersatti):

what is ur value of v=?

OpenStudy (anonymous):

v=5V

OpenStudy (qaisersatti):

ok wait

OpenStudy (qaisersatti):

what is the value of R in u r part 1

OpenStudy (anonymous):

833.333333333

OpenStudy (qaisersatti):

ok then do this 1.6vx-vout=0 is your first equation 100vx+ vout=500 is your second equation add both will give u vx= 4.92 v i= 4.92/R=4.92/833.333= 0.005905512 this will be your answer please check yourself too.. tc and if correct click the best choice:) bye

OpenStudy (anonymous):

worked.. tks, bro

OpenStudy (anonymous):

I didn't do the H12P2.. did you already do?

OpenStudy (anonymous):

H12P2 a) vin*(R1+R2)/2 b) vz*(R1+R2)/2 c? d? e? f? Have many problems with check. (BUG IN SYSTEM)

OpenStudy (anonymous):

any one help H12P3 last question.................

OpenStudy (anonymous):

use d formula above given by machu or gimme ur question :)

OpenStudy (anonymous):

any one did lab 12 please help how to calculate R1 C1 R2 C2